Night report 2021-01-01

Night report 2021-01-01

Observers: Priscila Pessi, Ken Smith

Support: Claudia Gutiérrez

TAT: Lluís Galbany, Giuliano Pignata, Sheng Yang, Janet Chen

EFOSC2 Calibrations:

Spectroscopic arcs

Biases

Domeflats

skyflats (there were a lot of issues with these flats, I don't know if they are usable)

SOFI calibrations:

JHK Flats

Weather:

Start: Humidity 48%, wind 1.2 m/s

Middle: humidity 47% wind 3.1 m/s

End: humidity 50% wind 5.9 m/s

Comments:

EFOSC2:

UT 01:20 image analysis. There were UPs (scientific term to define unknown problems) here with the instrument while doing image analysis. We spent some time here.

In the end we couldn't do image analysis. In addition the telescope was having problems on one axis.

We also had to do focus again. And it failed, so we try observing without focus and the telescope failed again.

Apparently there was an issue with the telescope during the afternoon and it was not completely fixed.

The telescope kept raising errors, I got bored so I kept describing them.

So far the night goes: pick object, try to do focus, fail. Try not to do focus, get an error. Change the object for another at lower airmass and repeat.

Now M1 can not be initialized.

UT03:10. success. Although, no image analysis was made, just focus.

While the telescope was not working the seeing was 0.7", when observing it was > 1.2" and it reached 2.3" at some point.

UT04:31 the altitude axis of the telescope crashed again while trying to get a standard at aimass 1.1. This time the recovery was made quickly.

Switch to SOFI

Č